# K-Space-Aware Cross-Modality Score for Synthesized Neuroimage Quality Assessment

### v1

@@ Abstract
The problem of how to assess cross-modality medical image synthesis has been largely unexplored. The most used measures like PSNR and SSIM focus on analyzing the structural features but neglect the crucial lesion location and fundamental k-space speciality of medical images. To overcome this problem, we propose a new metric K-CROSS to spur progress on this challenging problem. Specifically, K-CROSS uses a pre-trained multi-modality segmentation network to predict the lesion location, together with a tumor encoder for representing features, such as texture details and brightness intensities. To further reflect the frequency-specific information from the magnetic resonance imaging principles, both k-space features and vision features are obtained and employed in our comprehensive encoders with a frequency reconstruction penalty. The structure-shared encoders are designed and constrained with a similarity loss to capture the intrinsic common structural information for both modalities. As a consequence, the features learned from lesion regions, k-space, and anatomical structures are all captured, which serve as our quality evaluators. We evaluate the performance by constructing a large-scale cross-modality neuroimaging perceptual similarity (NIRPS) dataset with 6,000 radiologist judgments. Extensive experiments demonstrate that the proposed method outperforms other metrics, especially in comparison with the radiologists on NIRPS.
Index Terms:
Medical image, quality assessment, synthesized neuroimages, k-space
I Introduction
PSNR [15], SSIM [35], and MAE [7] are the most commonly used evaluation metrics in cross-modality magnetic resonance imaging (MRI) synthesis works. However, these metrics are inappropriate to a certain degree, considering that they are based on natural images and naturally ignore the inherent properties of MRI data. In general, the quality of neuroimage can be assessed by the content (i.e., lesion region), frequency space, and structure details. Although MAE, PSNR and SSIM are effective in assessing image quality, they are ineffective as a neuroimage metric, because they only focus on the structural details in the pixel space. Therefore, it is important to find a new way to measure how good the cross-modality neuroimage synthesis is.
@@ I Introduction
Besides, PSNR and SSIM are unable to account for differences in the k-space between the synthesized images and the target modality data, whereas K-CROSS can. The fundamental difference between MRI and natural images, as seen from the standpoint of imaging principles, is the basis of MR image reconstruction. The Fourier transformation, often known as the ”k-space” in MRI, is a mathematical concept that calculates various frequencies mixed into the received signal of all spins. It forms the basis for all image reconstruction in MRI. Therefore, we believe that the proposed metric can estimate the distance between MRIs in both k-space and pixel space. In k-space, the sophisticated K-CROSS encoder can capture the invariant modal-specific feature, where the frequency loss can be used to further enhance the complicated encoder. When a k-space shift occurs, as seen at the bottom of Fig. 1 (B), K-CROSS is more stable in accordance with the radiologist’s score, which is able to measure the gap in k-space between the synthesized neuroimage and the corresponding ground truth.
@@
The rest of this paper is organized as follows: Section II presents a literature review on image quality assessment and GAN-based assessment methods. Section III explains the proposed algorithm K-CROSS in detail. In addition, a large-scale multi-modal neuroimaging perceptual similarity (NIRPS) dataset is constructed in Section IV. Section V presents comprehensive experimental evaluations while Section V draws the conclusion and limitation of the current work.
II Related Work
II-A Image Quality Assessment
@@
Image quality assessment (IQA) can be divided into two categories. One is fully referenced IQA, and the other is non-referenced IQA [24, 25, 5, 33]. IQA with all references refers to estimating the quality of natural images with references: SSIM [35], MS-SSIM [36] and FSIM [40], focus more on image structure specifics. Specifically, FSIM builds up a novel feature similarity index according to the phase congruence and image gradient magnitude, while PSNR focuses on edge estimation for the synthesized images. Most of them [20, 26, 32, 42] use low-level features for evaluation. LPIPS [41] is the first work that uses a high-level feature for fully referenced IQA in light of the popularity of deep learning. Estimating the synthesized image quality without a reference (ground truth) is known as non-referenced IQA [24, 25, 5, 33]. RankIQA [22] is the mainstream for non-referenced IQA. Considering the limited size of IQA, Liu et al. propose a Siamese network to rank images and their distorted ones. The Siamese network’s knowledge (ranking result) can be transferred to a conventional neural network, whose function is to assess the quality of a single image. Since K-CROSS requires a reference image for evaluation, it belongs to a fully referenced IQA. However, few public data in the medical imaging community could be used to train for the learning-based fully referenced IQA methods. The NIRPS dataset, the first extensive neuroimaging perceptual similarity dataset with radiologists’ labels is constructed. As for fully referenced IQA methods, K-CROSS is, therefore, able to use the supervised training methods.
Figure 2: Flowchart of our proposed K-CROSS. First stage input: For reference neuroimage (source modality) and its query neuroimage (the target modality), the private portion is indicated by a blue box. The neuroimage’s structural feature is captured by the structure encoder. The reconstruction specifics are evaluated using $L_{stru}$. With a similarity loss $L_{simi}$, the presentation of the shared structure is maintained. The k-space feature to a modality is captured by Complex U-Net and optimized by $L_{freq}$. SegNet gets the mask of the tumor region from neuroimages that are specific to a given modality. The parameters of the off-the-shelf SegNet are not updated during the training phase, such as nnUnet [16], TransUNet [9], and SwinUNet [6]. The tumor encoder learns how to represent the tumor mask region, particularly texture details and the level of brightness. The quality of the SegNet-reconstructed tumor region is constrained using the loss $L_{tumor}$. First stage output: The private tumor encoder, the private complex encoder and the shared structure encoder for both modalities. Second stage input: The input are the query modality neuroimage and the modality-specific tumor encoder, complex encoder and the shared structure encoder from the first stage. The two main components of K-CROSS are $\eta_{complex}$ and $\eta_{nature}$. The complex score network yields $\eta_{complex}$, whereas the natural score network yields $\eta_{nature}$. The output of the tumor encoder and the structure encoder are combined as the input of the natural score network. The average score for $\eta_{complex}$ and $\eta_{nature}$ is $\eta_{total}$. For $\eta_{total}$, K-CROSS uses a straightforward regression model during the training phase, with labels taken from the NIPRS dataset. Inference: The input are the query modality neuroimage and the modality-specific tumor encoder, complex encoder and the shared structure encoder from the first stage. The output score is $\eta_{total}$.
II-B GAN Assessment
@@
The existing sample-based methods [38] have been proposed to access GAN performance, like Kernel MMD [12], Inception Score [27], Mode Score [8] and FID [13]. The classical approach is to compare the log-likelhood of generative models. But this approach cannnot accurately indicate the quality of synthesized image. In other words, a model can achieve high likelihood, but low image quality, and conversely, low image quality, and conversely. As for Inception score [27], it computes the KL divergence between the conditional class distribution and the marginal class distribution over the generated data. However, IS does not capture intra-class diversity, which is insensitive to the prior distribution over labels. Among them, the most popular metric is FID. Heusel et al. [13] use InceptionV3 [31] to extract the features from the real and synthetic neuroimaging data, and then compute the differences in the features between them. However, the majority of them are created in the pixel space and ignore the lesion region and k-space, which are the fundamental elements of MR image properties. In this regard, K-CROSS considers the underlying MR imaging principle as well as the difference between the neuroimages of the source and target modality.
III Proposed Method
III-A Preliminary
III-A1 K-Space Representations
@@
Figure 3: The architecture of complex encoder and decoder. The encoder consists of Complex Conv2d in Section III-A2 and Complex BatchNorm in Section III-A6. The decoder contains ComplexConvTranspose2d and Complex BatchNorm. The ComplexConvTranspose2d is similar to ComplxeConv2d except for the convolution operator. The middle part includes Complex Upsample in Section III-A7, Complex Tanh in Section III-A5 and Complex Conv2d.
@@
III-A2 Complex Convolution
The complex-valued convolution [34] is different from the real-valued convolution. Given a complex-valued convolution filter $\textbf{{W}}=\textbf{{A}}+i\textbf{{B}}$ with real-valued matrices A and B. The operation is expressed as follows:
@@ III-A2 Complex Convolution
III-A3 Complex Leaky RELU
It applies separate Leaky RELUs [23] on both the real part R(z) and the imaginary part Im(z) of a complex-valued, which is defined as:
@@ III-A3 Complex Leaky RELU
III-A4 Complex RELU
It applies separate RELUs [1] on both the real part R(z) and the imaginary part Im(z) of a complex-valued, which is defined as:
@@ III-A4 Complex RELU
III-A5 Complex Tanh
Complex Tanh applies separate tanh activation [18] on both the real part R(z) and imaginary part Im(z) of a complex-valued, which is defined as:
@@ III-A5 Complex Tanh
III-A6 Complex BatchNorm
As described in Cogswell et al. [10], complex-valued batch normalization could be separately applied into the imaginary part and real part, which could reduce the risk of over-fitting. The detail operation is defined as:
@@ III-A7 Complex Upsample
III-B Architecture
III-B1 Complex Branch
@@
III-B2 Tumor and Structure Branch
As a cross-modality segmentation neural network, the well-trained nnU-Net [16] is used in K-CROSS, with the weights being adjusted in the second stage of training. The modality-specific tumor encoder and decoder are private because the tumor information (the texture details and brightness) from the source modality and the target modality differ. With the exception of the operators using the normal convolution, batch norm, and Leaky RELU, the architecture details of the tumor encoder-decoder and the structure encoder-decoder are similar to those of Fig. 3.
III-B3 Score Network and Quality Prediction Regressor
@@
III-C Loss Function
III-C1 Frequency Loss
@@
Our solution is based on the focal frequency loss [17], as shown in Fig. 7. The hidden k-space of real MRI is $F_{r}(u,v)=a_{r}+b_{r}i$, and the corresponding k-space of synthesis MRI is $F_{f}(u,v)=a_{f}+b_{f}i$. To calculate their distance, we map $F_{r}$ and $F_{f}$ into the Euclidean space as $\vec{v_{r}}$ and $\vec{v_{f}}$. Specifically, the lengths of $\vec{v_{r}}$ and $\vec{v_{f}}$ are the amplitudes of $F_{r}$ and $F_{f}$, respectively. And the angles $\theta_{r}$ and $\theta_{f}$ correspond to the phases of $F_{r}$ and $F_{f}$, respectively. As a result, the distance between $F_{r}$ and $F_{f}$ can be converted to the distance between $\vec{v_{r}}$ and $\vec{v_{f}}$ (termed as $d(\vec{v_{r}},\vec{v_{f}})$), which is defined as follows:
@@
III-C2 Similarity Loss
For the similarity loss $\mathcal{L}_{simi}$, K-CROSS uses the maximum mean discrepancy (MMD) loss [30] to measure it. That is, K-CROSS computes the squared population MMD between shared structure encoding of the source modality $h^{s}_{c}$ and the target modality $h^{s}_{t}$ using a biased statistic. We express this as:

### v2

@@ Abstract
The problem of how to assess cross-modality medical image synthesis has been largely unexplored. The most used measures like PSNR and SSIM focus on analyzing the structural features but neglect the crucial lesion location and fundamental k-space speciality of medical images. To overcome this problem, we propose a new metric K-CROSS to spur progress on this challenging problem. Specifically, K-CROSS uses a pre-trained multi-modality segmentation network to predict the lesion location, together with a tumor encoder for representing features, such as texture details and brightness intensities. To further reflect the frequency-specific information from the magnetic resonance imaging principles, both k-space features and vision features are obtained and employed in our comprehensive encoders with a frequency reconstruction penalty. The structure-shared encoders are designed and constrained with a similarity loss to capture the intrinsic common structural information for both modalities. As a consequence, the features learned from lesion regions, k-space, and anatomical structures are all captured, which serve as our quality evaluators. We evaluate the performance by constructing a large-scale cross-modality neuroimaging perceptual similarity (NIRPS) dataset with 6,000 radiologist judgments. Extensive experiments demonstrate that the proposed method outperforms other metrics, especially in comparison with the radiologists on NIRPS. To forster reproducibility and accessibility, the source codeis uploaded to the website: https://github.com/M-3LAB/K-CROSS
keywords:
[…]
{highlights}
1 Introduction
PSNR (Huynh-Thu and Ghanbari, 2008), SSIM (Wang et al., 2004), and MAE (Chai and Draxler, 2014) are the most commonly used evaluation metrics in cross-modality magnetic resonance imaging (MRI) synthesis works. However, these metrics are inappropriate to a certain degree, considering that they are based on natural images and naturally ignore the inherent properties of MRI data. In general, the quality of neuroimage can be assessed by the content (i.e., lesion region), frequency space, and structure details. Although MAE, PSNR and SSIM are effective in assessing image quality, they are ineffective as a neuroimage metric, because they only focus on the structural details in the pixel space. Therefore, it is important to find a new way to measure how good the cross-modality neuroimage synthesis is. To forster reproducibility and accessibility, the source code is uploaded to .
@@ 1 Introduction
Besides, PSNR and SSIM are unable to account for differences in the k-space between the synthesized images and the target modality data, whereas K-CROSS can. The fundamental difference between MRI and natural images, as seen from the standpoint of imaging principles, is the basis of MR image reconstruction. The Fourier transformation, often known as the "k-space" in MRI, is a mathematical concept that calculates various frequencies mixed into the received signal of all spins. It forms the basis for all image reconstruction in MRI. Therefore, we believe that the proposed metric can estimate the distance between MRIs in both k-space and pixel space. In k-space, the sophisticated K-CROSS encoder can capture the invariant modal-specific feature, where the frequency loss can be used to further enhance the complicated encoder. When a k-space shift occurs, as seen at the bottom of Fig. 1 (B), K-CROSS is more stable in accordance with the radiologist’s score, which is able to measure the gap in k-space between the synthesized neuroimage and the corresponding ground truth.
@@
The rest of this paper is organized as follows: Section 2 presents a literature review on image quality assessment and GAN-based assessment methods. Section 3 explains the proposed algorithm K-CROSS in detail. In addition, a large-scale multi-modal neuroimaging perceptual similarity (NIRPS) dataset is constructed in Section 4. Section 5 presents comprehensive experimental evaluations while Section V draws the conclusion and limitation of the current work.
2 Related Work
2.1 Image Quality Assessment
@@
Image quality assessment (IQA) can be divided into two categories. One is fully referenced IQA, and the other is non-referenced IQA (Mittal et al., 2013; Moorthy and Bovik, 2011; Bosse et al., 2016; Talebi and Milanfar, 2018). IQA with all references refers to estimating the quality of natural images with references: SSIM (Wang et al., 2004), MS-SSIM (Wang et al., 2003) and FSIM (Zhang et al., 2011), focus more on image structure specifics. Specifically, FSIM builds up a novel feature similarity index according to the phase congruence and image gradient magnitude, while PSNR focuses on edge estimation for the synthesized images. Most of them (Lin and Wang, 2018; Ren et al., 2018; taek Lim et al., 2018; Zhang et al., 2021) use low-level features for evaluation. LPIPS (Zhang et al., 2018) is the first work that uses a high-level feature for fully referenced IQA in light of the popularity of deep learning. Estimating the synthesized image quality without a reference (ground truth) is known as non-referenced IQA (Mittal et al., 2013; Moorthy and Bovik, 2011; Bosse et al., 2016; Talebi and Milanfar, 2018). RankIQA (Liu et al., 2017b) is the mainstream for non-referenced IQA. Considering the limited size of IQA, Liu et al. propose a Siamese network to rank images and their distorted ones. The Siamese network’s knowledge (ranking result) can be transferred to a conventional neural network, whose function is to assess the quality of a single image. Since K-CROSS requires a reference image for evaluation, it belongs to a fully referenced IQA. However, few public data in the medical imaging community could be used to train for the learning-based fully referenced IQA methods. The NIRPS dataset, the first extensive neuroimaging perceptual similarity dataset with radiologists’ labels is constructed. As for fully referenced IQA methods, K-CROSS is, therefore, able to use the supervised training methods.
2.2 GAN Assessment
@@
The existing sample-based methods (Xu et al., 2018) have been proposed to access GAN performance, like Kernel MMD (Gretton et al., 2012), Inception Score (Salimans et al., 2016), Mode Score (Che et al., 2017) and FID (Heusel et al., 2017). The classical approach is to compare the log-likelhood of generative models. But this approach cannnot accurately indicate the quality of synthesized image. In other words, a model can achieve high likelihood, but low image quality, and conversely, low image quality, and conversely. As for Inception score (Salimans et al., 2016), it computes the KL divergence between the conditional class distribution and the marginal class distribution over the generated data. However, IS does not capture intra-class diversity, which is insensitive to the prior distribution over labels. Among them, the most popular metric is FID. Heusel et al. (Heusel et al., 2017) use InceptionV3 (Szegedy et al., 2016) to extract the features from the real and synthetic neuroimaging data, and then compute the differences in the features between them. However, the majority of them are created in the pixel space and ignore the lesion region and k-space, which are the fundamental elements of MR image properties. In this regard, K-CROSS considers the underlying MR imaging principle as well as the difference between the neuroimages of the source and target modality.
2.3 Cross-Modality Medical Image Synthesis
Existing medical image-to-image translation (Jiang et al., 2019; Ren et al., 2021; Kong et al., 2021) has demonstrated their considerable research and clinical analysis potential. Of these methods, supervised GANs are still the mainstream for cross-modality neuroimaging data synthesis (Wang et al., 2018; Dar et al., 2019; Sharma and Hamarneh, 2020; Yu et al., 2020, 2018; Zuo et al., 2021). For instance, Elad et al. (Elad, 2012) provide a concise overview of sparse and redundant representation modeling and identify ten important future research directions for sparse coding. Rubinstein et al. (Rubinstein et al., 2010) explicate the dictionary acquisition process through mathematical models. Gao et al. (Gao et al., 2012) develop a hypergraph Laplacian matrix to retain the local information of the training samples, thereby enhancing the discriminative capacity of the learned dictionary. However, synthesizing in a supervised manner requires paired data for training, which is difficult to implement in practice. To solve this problem, both semi-supervised and unsupervised methods are then launched to eliminate the need of paired data. (Guo et al., 2021) leverage a lesion segmentation network as a teacher to guide the generator by using unpaired training data. (Shen et al., 2021) and (Zhou et al., 2021) also utilize the high-level tasks to guide the cross-modality image synthesis. Huang et al. (Huang et al., 2020b, a) make full use of unpaired cross-modality data and project them into a common space. The attributed features from the common space bring great helpful to synthesize the missing target modality data. (Li et al., 2022) employ the dual-domain attention mechanism to extract highly discriminative features on the lesion area. Yu et al. (Yu et al., 2021) present a similar work with the method shown in (Huang et al., 2020a). However, the authors pay more attention to the mouse brain dataset. Chen et al. (Chen et al., 2020) propose a more concise idea, i.e., the source modality and the target modality share their feature encoder. Jiao et al. (Jiao et al., 2020) also extract and map features into the common space using different modalities. Moreover, the authors in (Jiao et al., 2020) design a new cross-modal attention module for fusion and propagation. Zeng et al. (Zeng and Zheng, 2019) use two models; one of which is the 3D generator network, and the other is the 2D discriminator. The authors utilize the result from the 2D discriminator treated as a weak label to supervise the 3D generator, such that the generator’s output can be closer to the output of CT. Yang et al. (Yang et al., 2021) design a unified generator for MRI synthesis. (Wu et al., 2023) propose a MRI oriented novel attention-based glioma grading network into mutli-scale feature extraction process, which aims to promote the synergistic interaction among different modality information. (Kong et al., 2021) introduce a new I2IT model called RegGAN, which converts the unsupervised I2IT task into a supervised I2IT with noisy labels. K-CROSS serves as a general metrics for various levels of supervision of cross-modality medical image synthesis.
Figure 2: Flowchart of our proposed K-CROSS. First stage input: For reference neuroimage (source modality) and its query neuroimage (the target modality), the private portion is indicated by a blue box. The neuroimage’s structural feature is captured by the structure encoder. The reconstruction specifics are evaluated using $L_{stru}$. With a similarity loss $L_{simi}$, the presentation of the shared structure is maintained. The k-space feature to a modality is captured by Complex U-Net and optimized by $L_{freq}$. SegNet gets the mask of the tumor region from neuroimages that are specific to a given modality. The parameters of the off-the-shelf SegNet are not updated during the training phase, such as nnUnet (Isensee et al., 2020), TransUNet (Chen et al., 2021), and SwinUNet (Cao et al., 2021). The tumor encoder learns how to represent the tumor mask region, particularly texture details and the level of brightness. The quality of the SegNet-reconstructed tumor region is constrained using the loss $L_{tumor}$. First stage output: The private tumor encoder, the private complex encoder and the shared structure encoder for both modalities. Second stage input: The input are the query modality neuroimage and the modality-specific tumor encoder, complex encoder and the shared structure encoder from the first stage. The two main components of K-CROSS are $\eta_{complex}$ and $\eta_{nature}$. The complex score network yields $\eta_{complex}$, whereas the natural score network yields $\eta_{nature}$. The output of the tumor encoder and the structure encoder are combined as the input of the natural score network. The average score for $\eta_{complex}$ and $\eta_{nature}$ is $\eta_{total}$. For $\eta_{total}$, K-CROSS uses a straightforward regression model during the training phase, with labels taken from the NIPRS dataset. Inference: The input are the query modality neuroimage and the modality-specific tumor encoder, complex encoder and the shared structure encoder from the first stage. The output score is $\eta_{total}$.
3 Proposed Method
3.1 Preliminary
3.1.1 K-Space Representations
@@
Figure 3: The architecture of complex encoder and decoder. The encoder consists of Complex Conv2d in Section 3.1.2 and Complex BatchNorm in Section 3.1.6. The decoder contains ComplexConvTranspose2d and Complex BatchNorm. The ComplexConvTranspose2d is similar to ComplxeConv2d except for the convolution operator. The middle part includes Complex Upsample in Section 3.1.7, Complex Tanh in Section 3.1.5 and Complex Conv2d.
@@
3.1.2 Complex Convolution
The complex-valued convolution (Trabelsi et al., 2018) is different from the real-valued convolution. Given a complex-valued convolution filter $\textbf{{W}}=\textbf{{A}}+i\textbf{{B}}$ with real-valued matrices A and B. The operation is expressed as follows:
@@ 3.1.2 Complex Convolution
3.1.3 Complex Leaky RELU
It applies separate Leaky RELUs (Maas, 2013) on both the real part R(z) and the imaginary part Im(z) of a complex-valued, which is defined as:
@@ 3.1.3 Complex Leaky RELU
3.1.4 Complex RELU
It applies separate RELUs (Agarap, 2018) on both the real part R(z) and the imaginary part Im(z) of a complex-valued, which is defined as:
@@ 3.1.4 Complex RELU
3.1.5 Complex Tanh
Complex Tanh applies separate tanh activation (Kalman and Kwasny, 1992) on both the real part R(z) and imaginary part Im(z) of a complex-valued, which is defined as:
@@ 3.1.5 Complex Tanh
3.1.6 Complex BatchNorm
As described in Cogswell et al. (Cogswell et al., 2016), complex-valued batch normalization could be separately applied into the imaginary part and real part, which could reduce the risk of over-fitting. The detail operation is defined as:
@@ 3.1.7 Complex Upsample
3.2 Architecture
The two stages of training for K-CROSS are depicted in Fig 2. Encoder-decoder architecture is primarily used in the first stage of K-CROSS to train the tumor encoder, complex encoder, and structure encoder. The complex score network and natural score network are optimized by K-CROSS using a regressor in the second stage. For stage one, the source modality neuroimage and the target modality neuroimage are fed into tumor branch, complex branch and structure branch. A shared cross-modality segmentation network and a personal tumor encoder-decoder make up the tumor branch. The segmentation network is used to obtain the predicted lesion region. The tumor encoder then captures the brightness and texture features of the lesion region. As for complex branch, the input is converted into k-space by Fourier transform. The function of the private complex encoder is to identify the k-space feature shit. The source modality and the target modality share the structure branches in K-CROSS. The structure’s goal is to gather information about the shared structure from both inputs. The architecture and weights of the tumor encoder, complex encoder, and structure encoder are cloned from the first stage in the second stage. The complex score network and the natural score network are optimized by K-CROSS using a regressor.
3.2.1 Complex Branch
@@
3.2.2 Tumor and Structure Branch
As a cross-modality segmentation neural network, the well-trained nnU-Net (Isensee et al., 2020) is used in K-CROSS, with the weights being adjusted in the second stage of training. The modality-specific tumor encoder and decoder are private because the tumor information (the texture details and brightness) from the source modality and the target modality differ. With the exception of the operators using the normal convolution, batch norm, and Leaky RELU, the architecture details of the tumor encoder-decoder and the structure encoder-decoder are similar to those of Fig. 3.
3.2.3 Score Network and Quality Prediction Regressor
@@
3.3 Loss Function
3.3.1 Frequency Loss
@@
Our solution is based on the focal frequency loss (Jiang et al., 2021), as shown in Fig. 7. The hidden k-space of real MRI is $F_{r}(u,v)=a_{r}+b_{r}i$, and the corresponding k-space of synthesis MRI is $F_{f}(u,v)=a_{f}+b_{f}i$. To calculate their distance, we map $F_{r}$ and $F_{f}$ into the Euclidean space as $\vec{v_{r}}$ and $\vec{v_{f}}$. Specifically, the lengths of $\vec{v_{r}}$ and $\vec{v_{f}}$ are the amplitudes of $F_{r}$ and $F_{f}$, respectively. And the angles $\theta_{r}$ and $\theta_{f}$ correspond to the phases of $F_{r}$ and $F_{f}$, respectively. As a result, the distance between $F_{r}$ and $F_{f}$ can be converted to the distance between $\vec{v_{r}}$ and $\vec{v_{f}}$ (termed as $d(\vec{v_{r}},\vec{v_{f}})$), which is defined as follows:
@@
3.3.2 Similarity Loss
For the similarity loss $\mathcal{L}_{simi}$, K-CROSS uses the maximum mean discrepancy (MMD) loss (Simonyan and Zisserman, 2015) to measure it. That is, K-CROSS computes the squared population MMD between shared structure encoding of the source modality $h^{s}_{c}$ and the target modality $h^{s}_{t}$ using a biased statistic. We express this as: